OpenStudy (anonymous):

Can 3( 1x^2 + 7x + 10) be factored completely?

OpenStudy (anonymous):

=3(x+5)(x+2) since 5*2=10 and 5+2=7

OpenStudy (anonymous):

OOOO :O thanks man

OpenStudy (anonymous):

ur welcome :})